The setting of cash collection targets is important to any credit department in order to control working capital efficiently. Targets also act as a motivational tool, as without them collection activity is likely to be aimless and results, at best, mediocre. It is, therefore, essential that a credit manager knows how to forecast accounts receivable levels in both debtor days and monetary values.
The following figures are available to you:
£’000 £’000
Accounts receivable as at 30 June 4,550
Sales:   
June 2,100   
May 2,250   
April 2,020   
Budgeted Sales:   
July 2,167   
August 2,000   
September 2,200   
You are required to:
i) Calculate the Days Sales Outstanding (DSO) as at 30 June, using both the adding back and annualised methods. Show your calculations and briefly comment on the results.
ii) Using 30 days for each month and showing the methods used, calculate the amount of money the department would need to collect to achieve forecasted DSO figures of 61 days in July.
iii) Discuss additional measures that can be used to monitor the performance of the collections department.
iv) Describe how the performance of the department could be reported to the Board of Directors at month end.
Answer #1

i) Account receivables as at 30th june = 4550

sales value on june = 2100 (it is taken as credit sale)

In june there are 30 days, so the number of days = 30

DSO = Account Receivables / Net credit sales * Number of days

= 4550 / 2100 * 30 = 65

so here, the DSO will be 65. (Generally the DSO under 45 days is considered as excellent). To conclude that the DSO obtained here is not satisfactory.

ii) DSO calculation for month MAY

credit sales in the month May = 2250

assume that the account receivables in May = 4700 ( 2250 - 2100 = 150, adding the 150 to 4550)

Number of days = 30

DSO = 4700 / 2250 * 30 = 62.667 ( so here the DSO is not favourable)

DSO FOR APRIL

Credit sales = 2020

Account receivables in April = 4470 (2100 - 2020 = 80, less the 80 from 4550)

Number of days = 30

DSO = 4470 / 2020 * 30 = 66.38, ( to conclude that the DSO obtained here is not favourable)

DSO FOR JULY

Credit sales = 2167

Account receivable = 4617 ( 2167 - 2100 = 67, add the 67 to 4550)

Number of days = 30

DSO = 4617 / 2167 * 30 = 63.9 or 64. ( also here the DSO is not favourable)

IF the DSO is 61 in july, then the receivable :

DSO = Account receivables / credit sales * number of days

here the ; DSO = 61, credit sales = 2167 , Number of days = 30

61 = Account receivables / 2167 * 30

Account Receivables * 30 = 61 * 2167

Account receivables * 30 = 132187

therefore the account receivables = 132187 / 30 = 4406.23 or 4406.

DSO FOR AUGUST

Credit sale = 2000

Account receivables = 4450 ( 2100 - 2000 = 100, less the 100 from 4550)

Number of days = 30

DSO = 4450 / 2000 * 30 = 66.75, ( the DSO is not favour in position)

DSO FOR SEPTEMBER

Credit sale = 2200

Account recivables = 4650 ( 2200 - 2100 = 100, add the 100 to 4550)

Number of days = 30

DSO = 4650 / 2200 * 30 = 63.40, ( so here the DSO is not favourable)

iii) There are other different methods are available for the monitoring the performance of collection departments;

1) Collector Effective Index (CEI) ; it also measure the amount collected for the specific period of time against the total receivable during the same time. this tool measure the performance in percentage ( 100%). the tool also helps to plan for the future.

2) Right Party Contact (RCP) ; it measures the ratio of all outgoing calls made to the valid phone number of the person from whom collection is sought. higher the ratio indicate the high access of debt collection.

3) Percentage of outbound calls resulting the promise to pay (PTP);   this is a metrix that measures success rate of the RCP contacts. it also measure all the calls that eded with promise to pay. this tool also can be express in the 100 per cent.

4) Profit Per Account (PPA) ; this is another tool that measures that the profit generated from the average debt collection during a particular period. this tool also helps to analyse the impact of various inputs in account receivables.

iv) From the above data analysed, it is clear that the companies cash collection target is very ineffective, the reason is that ; every months debt collection efficiency ie, the DSO is above 60. which is not much favourable for the firm. generally the DSO below 45 is considered as excellent. Depending upon the credit sale and the account receivable the account receivable is huge and sales are minimum in nature. In the month April DSO is 66.38 , May - 62.67, June - 65, July - 64, August - 66.75, September - 63.40. In every month DSO will fluctuate. The company need to increase its debt collection speed and reduce its account receivables.